OpenStudy (anonymous):

FINAL REVIEW HELP- use the product rule to simplify the radical : sqrt(150)

OpenStudy (raden):

sqrt(150) = sqrt(25*6) = sqrt(25) * sqrt(6) right ?

OpenStudy (chihiroasleaf):

\[\sqrt{ab} = \sqrt{a} \times \sqrt{b}\]

OpenStudy (anonymous):

is the answer\[5\sqrt{6}\]

OpenStudy (raden):

yeah, that is a start

OpenStudy (anonymous):

?

OpenStudy (raden):

u are right ... sorry , my conection was lagging :(

OpenStudy (anonymous):

Its ok, Thanks!

OpenStudy (raden):

welcome